OpenStudy (anonymous):

Find two positive real numbers whose sum is 100 and whose product is a maximum?

OpenStudy (whpalmer4):

You can think of this as \[y = x*(100-x)\] which is an inverted parabola. What is the vertex?

OpenStudy (anonymous):

Domain is (100, infinity)?

OpenStudy (ash2326):

@kissy Let the numbers be x and y we have a condition \[x+y=100\] Do you get this part?

OpenStudy (anonymous):

absolutely

OpenStudy (ash2326):

Have you studies calculus?

OpenStudy (ash2326):

*studied

OpenStudy (ash2326):

I just want to know if you know differentiation and how to find the maxima and minima for a function. Of course I'll explain

OpenStudy (anonymous):

Oh yes, I understand how to differentiate and how to find the extrema. I'm just having trouble understanding how to solve optimization problems

OpenStudy (ash2326):

Ok, let's work on it

OpenStudy (ash2326):

Let the product of x and y be P so \[P= x\times y\] We know y=100-x so \[P=x\times (100-x)\] Do you get this part?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

So we want to maximize P Let's differentiate P with respect to x. Can you do that?

OpenStudy (anonymous):

use chain rule to differentiate, right?

OpenStudy (anonymous):

so P'(x)= 1(100-x)*(1)

OpenStudy (ash2326):

It should be \[P'(x)=1 \times (100-x)+x\times (0-1)\]

OpenStudy (anonymous):

Shoot, forgot the x. Sorry

OpenStudy (ash2326):

No problem, so we get \[P'(x)=100-x-x\] Can you find value of x is P'(x)=0

OpenStudy (anonymous):

10?

OpenStudy (ash2326):

\[0=100-x-x\] \[0=100-2x\] Can you try now?

OpenStudy (anonymous):

Oh, I thought you meant what makes P'(x) 0 but you meant make P'(x) equal to 0. It is 50

OpenStudy (ash2326):

Cool. I understood that. So there you go you got x , can you find y?

OpenStudy (anonymous):

To find y you would make P'(0) right? ... as in you plug 0 into the x's of the original formula

OpenStudy (ash2326):

We have the original condition \[x+y=100\] put x=50 and find y from this.

OpenStudy (anonymous):

50 then.... where did you get x+y=100?

OpenStudy (anonymous):

...nevermind. The questions states that the sum of x and y is 100 lol

OpenStudy (ash2326):

Cool, so these are the two no.s which will make product the maximum

OpenStudy (anonymous):

they are the CN's?

OpenStudy (ash2326):

CN?

OpenStudy (anonymous):

critical numbers

OpenStudy (ash2326):

yes

OpenStudy (ash2326):

do you understand?

OpenStudy (anonymous):

Yes I do. I'm still a bit confused still though... like I understand what you did and i'm just still not understanding the rest of the steps

OpenStudy (ash2326):

Just read it again, all of my posts. Let me know wherever you have doubt

OpenStudy (anonymous):

Alright. I will look through again and try to figure it out on my own but if I still don't understand it I will let you know (: thanks

OpenStudy (ash2326):

Cool

OpenStudy (anonymous):

no calculus needed the sum is 100, the product is a max, it is \(50\times 50\)

OpenStudy (anonymous):

I still don't quite understand

OpenStudy (anonymous):

ok you have two numbers whose sum is 100 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so for example one could be 30 and the other 70 or one could be 25 and the other 75 or one could be 99 and the other 1 or even one could be 100 and the other one 0 in other words, you have two number \(x\) and \(y\) where \(x+y=100\) or if you prefer you have two numbers \(x\) and \(100-x\)

OpenStudy (anonymous):

Oh, so the x and y that we just found are the two positive real numbers

OpenStudy (anonymous):

yeah, just some real numbers now here is the thing: you cannot tell then apart by which i mean if you say one is \(x\) so the other is \(y=100-x\) or i say one is \(100-x\) and the other is \(x\) we both have the same thing, two numbers whose sum is 100

OpenStudy (anonymous):

since you cannot distinguish between the numbers, \(x\) and \(y\) or \(x\) and \(100-x\), i.e. since they are interchangeable, the procedure for finding the maximum product cannot favor one number over the other you say \(x+y=100\) and i say \(y+x=100\) we are both saying the same thing

OpenStudy (anonymous):

so the maximum must be when the two numbers are identical, because we cannot tell them apart

OpenStudy (anonymous):

you can check it with numbers and see what i mean you say the first number is \(30\) so the second number must be \(70\) and \(30\times 70=2100\) now if i interchange them, we get \(70\times 30=2100\) the same answer exactly now they will be biggest if we make both numbers the same, i.e. if \(x=y=50\) and \(50\times 50=2500\) is the biggest

OpenStudy (anonymous):

Ok, I understand. So it'd be 50x50 for max and the two positive real numbers are 50 and 50 because like you said about the identical thing

OpenStudy (anonymous):

if you don't like my logic which requires only common sense, we can also do it using algebra

OpenStudy (anonymous):

right, you cannot tell them apart, so this procedure of finding the max cannot favor one number over the other but we can still use algebra if this will make your teacher happy

OpenStudy (anonymous):

Alright, how would you do it the algebra way? (it's not required but i'd still like to understand it)

OpenStudy (anonymous):

call one number \(x\) so the other number must be \(100-x\) and their product is \(x(100-x)=100x-x^2\)

OpenStudy (anonymous):

the equation \[y=100x-x^2\] is a parabola that opens down the biggest it can be is at the vertex, and the first coordinate of the vertex of a parabola is \(-\frac{b}{2a}\) which in this case is \(-\frac{100}{2\times (-1)}=50\)

OpenStudy (anonymous):

therefore the maximum the product can be is if \(x=50\) which of course means the other number is \(50\) as well, and \(50\times 50=2500\)

OpenStudy (anonymous):

it is simple enough algebra, but common sense is even simpler how could it be possible that the maximum of the product could anything other than the number you would get if the two numbers were equal? why would one number be favored over the other? in other words, since \(x+y=100\) is symmetric in \(x\) and \(y\) by which i mean \(x+y=100\) is exactly the same as \(y+x=100\) it must be the case that the max is if \(x=y\)

OpenStudy (anonymous):

Thank you!